The cost analyst at Fitzroy Services has conducted an analysis of the support cost accounts at their Victoria Branch. Fitzroya Fixed cost for Supervision b. Total variable cost for Supervision c. Number of customer transactions
Answer #1

a) Fixed cost for supervision

Total fixed cost 120100
Supplies -17400
Customer support -39900
Administration -54400
Supervision 8400

b) Variable cost for supervision = 47000-8400 = 38600

c) Number of customer transaction =

230400 = 120100+0.20X

110300 = 0.20X

Number of customer = 110300/.20 = 551500 customers